OpenStudy (anonymous):

is the trig. ratio of cos300 degrees 1/2?

OpenStudy (anonymous):

yup cos300=cos(360-60)=cos60=1/2

OpenStudy (anonymous):

okay, so the tan315 would be 1?

zepdrix (zepdrix):

Hmm that would be -1 c: